# Leveraging Local Patch Alignment to Seam-cutting for Large Parallax Image Stitching

### v2

@@ Abstract
Seam cutting methods have been proven effective in the composition step of image stitching, especially for images with parallax. However, current seam cutting can be seen as the subsequent step after the image alignment is settled. Its effectiveness usually depends on the fact that images can be roughly aligned such that a local region exists where an unnoticeable seam can be found. Current alignment methods often fall short of expectations for images with large parallax, and most efforts are devoted to improving the alignment accuracy.
In this paper, we argue that by adding a simple Local Patch Alignment Module (LPAM) into the seam cutting, the final result can be efficiently improved for large parallax image stitching. Concretely, we first evaluate the quality of pixels along the estimated seam of the seam cutting method. Then, for pixels with low qualities, we separate their enclosing patches in the aligned images and locally align them by constructing modified dense correspondences via SIFT flow. Finally, we composite the aligned patches via seam cutting and merge them into the original aligned result to generate the final mosaic. Experiments show that introducing LPAM can effectively and efficiently improve the stitching results.
Index Terms:
Image stitching, large parallax, seam-cutting, local patch alignment, SIFT flow.
I Introduction
Image stitching has been well-studied for decades in computer vision [1], which mainly involves image alignment [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], composition and blending [15, 16, 17, 18, 19, 20, 21]. In consumer-level photography, parallax is the most intractable issue due to the unconstrained shooting environment. Existing alignment methods may still fail to handle it. Therefore, it is usually the image composition that determines the quality of the final stitching result.
seam cutting [18, 17, 19, 22, 20, 21] and seam-driven approaches [23, 24, 25], which aim to find an optimal seam in the overlapping region from one or finite alignment candidates, have been proven effective in handing parallax. However, as a second step after image alignment is settled in image stitching, the effectiveness highly depends on the alignment accuracy. It means that the images should at least be roughly aligned such that there exists a local region where an unnoticeable seam can be found. For images with large parallax, existing alignment methods may fail to align them locally (see Fig. 1(b)), resulting in the seam cutting failing to produce visually pleasing results (see Fig. 1(c,d)).
(a) Input images
(b) Averaged aligned images
(c) Method [19]
(d) Method [20]
(e) Our method
@@ I Introduction
Figure 1: Issues of the state-of-the-art seam cutting methods for large parallax image stitching (marked by red arrows). (b) Aligned images are obtained via method [6], where the overlapping region is averaged. The curbs are totally misaligned (Best view in color and zoom in).
@@
In this paper, we argue that by adding a simple Local Patch Alignment Module (LPAM) into the seam cutting, the final result can be efficiently improved for large parallax image stitching. Specifically, we first introduce a seam quality evaluation algorithm to evaluate the quality of pixels along the seam and identify the “problematic” pixels. Then, for these pixels, we separate their enclosing patches in the aligned images and locally align them by constructing modified dense correspondences via SIFT flow. Finally, we composite the aligned patches via seam cutting and merge them into the former aligned result to generate the final mosaic. Experiments show that adding our LPAM into the seam cutting can effectively and efficiently improve the stitching results (see Fig. 1(e)).
Our method has two major contributions.
  • We propose a local alignment and seam cutting method that can improve the performance of the large parallax image stitching methods.
  • Our method is a plug-and-play module that can be effectively and efficiently applied to any seam cutting method.
II Related Work
@@
II-A Seam cutting Methods
@@
The seam cutting approach is a powerful and widely used composition method [15, 28, 16, 17, 18, 19, 29], which is usually expressed in terms of energy minimization and minimize it via graph-cut optimization [30]. Agarwala et al. [28] applied seam cutting to generate “digital photomontage”, they defined the energy function as the sum of a data penalty selecting by the user and an interaction penalty involving the colors, gradients, and edges of an image.
@@
Zhang et al. [16] combined alignment error and color difference in the energy function to find a better seam that avoids splitting misaligned pixels with similar colors. Li and Liao [18] considered the nonlinearity and nonuniformity of human perception into seam cutting and defined a sigmoid-metric color difference and saliency weights in the energy function. To avoid objects from being cropped, omitted, or duplicated, Herrmann et al. [17] utilized object detection to match objects between images, and introduced three new terms to the traditional energy function that address the above issues, respectively. Liao et al. [19] indicated that the seam with minimal cost is not necessarily optimal in terms of human perception. Then, they proposed an iterative seam estimation method where a quality evaluation algorithm for pixels along the seam is designed to guide the iteration procedure. Dai et al. [22] proposed the first end-to-end deep learning framework for image composition. They cast the seam cutting as an image-blending problem and designed an edge-guided network to regress the blending weights and seamlessly produce the stitched image. Li and Zhou [20] utilize the quaternion representation of the color image to define the energy function and perform seam cutting completely in the quaternion domain.
II-B Seam-driven Methods
@@
The seam-driven strategy was first proposed by Gao et al. [23]. They indicated that the perceptually best result is not necessarily from the best global alignment. Thus, they generated finite alignment candidates from multiple RANSAC [31] procedures and applied seam cutting to estimate multiple seams from these candidates. Then, a seam quality metric is defined to evaluate these seams, and the seam with minimal cost is chosen to produce a final mosaic. Zhang and Liu [24] improved the screening process of alignment candidates by combining homography with content-preserving warps. They used the seam cost of the energy function to evaluate the seam quality. Lin et al. [25] generated the alignment candidates via a superpixel-based feature grouping and a structure-preserving warp. The warp is gradually improved via adaptive feature weighting. They also defined a seam quality metric based on the zero-mean normalized cross-correlation (ZNCC) score. Zhang et al. [21] proposed an iterative seam-based stitching method that leverages dense flow estimation to directly estimate seam in the input images and use it to solve a spatial smooth warp.
@@
In both seam cutting and seam-driven methods, their main focus is defining various energy functions to find an optimal seam from the settled pre-aligned images. For images with large parallax, these methods may easily fail if the alignment methods fail to locally align the images. Our method, on the other hand, improves the performance by realigning the local misaligned patches where the estimated seam passes.
Figure 2: An overview of the proposed local patch alignment module.
III Proposed architecture
@@
As depicted in Fig. 2, our architecture consists of two stages, seam quality evaluation and patch extraction, succeeded by patch alignment and seam merging. Subsequently, we first provide a preliminary of how seam cutting works, and then present a detailed description of the two stages.
III-A Preliminary
@@
Given aligned target image $I_{0}$ and reference image $I_{1}$, for the overlapping region $\mathbf{O}$, seam cutting method estimates an initial seam $S_{0}$ by finding a labeling $l$ that minimizes the energy function
@@
III-B Seam quality evaluation and patch extraction
Since the seam cutting methods mainly fail due to the structure misalignment, we use the structural similarity (SSIM) index [32] to evaluate each pixel $p_{i}$ on the seam $S_{0}$. Concretely, we extract two local patches $P_{0}$ and $P_{1}$ centered at $p_{i}$, and define the evaluation error for $p_{i}$ as
@@ III-B Seam quality evaluation and patch extraction
We set $k$ to 1.5 in our method. The threshold $\tau$ divides the seam into multiple components. Notably, the threshold might not be sufficiently precise to identify all misaligned pixels. However, we allow for some some false positives because the subsequent alignment process does not yield inferior results. For each misaligned component $\mathbf{C}_{i}$, we separate its enclosing rectangular patches $\mathbf{P}^{i}_{0}$ and $\mathbf{P}^{i}_{1}$ from the corresponding aligned images (see Fig. 2).
III-C Patch Alignment and seam merging
Figure 3: Illustration for patch alignment and seam merging process. Artifacts are denoted by red arrows. From left to right in (a,b,c,d): vector flow, local stitching seam, and local composited result. For clarity, only part of the vector flow is presented. From (a) to (d): results of original $V^{i}$, modified $f(V^{i})$ with $\beta=1,4,8$ in Eq. (5), respectively. The green rectangle in the middle of each subfigure indicates the position of the separated patch and the merging of the local seam. The right image in each subfigure only shows the simple composited results without gradient domain fusion (Best view in color and zoom in).
The separated patches exhibit structure misalignment, we assume that there is either no or only a small parallax between these patches. Thus, we can utilize SIFT flow [34] to construct dense correspondences among the pixels within the patches. Theoretically, for each pixel $p_{j}$ in $\mathbf{P}_{0}^{i}$, there exists a shift vector $V^{i}(p_{j})$ such that
@@ III-C Patch Alignment and seam merging
Then, the two patches can be aligned via the vector flow $V^{i}$. However, if the boundary pixels of the patch $\mathbf{P}_{0}^{i}$ have large shift vectors, merging the aligned patches into the pre-aligned results may introduce more artifacts around the patch, as shown in Fig. 3. To address this, we smoothly modify the vector flow and realign the two patches as
$\tilde{\mathbf{P}}_{0}^{i}(p_{j})=\mathbf{P}_{0}^{i}(p_{j}+f(V^{i}(p_{j}))),\quad\tilde{\mathbf{P}}_{1}^{i}(p_{j})=\mathbf{P}_{1}^{i}(p_{j})$ (4)
@@
For horizontal patch alignment, $t$ represents the $x$-coordinate of a pixel within patch $\mathbf{P}_{0}^{i}$, normalized to $[0,1]$. $\beta$ controls the rate of change of the function $f$. The sigmoid function ensures that the aligned patches are smoothly consistent with $I_{0}$ when $t=0$ and with $I_{1}$ when $t=1$. As shown in Fig. 3, the modified vector flow under different settings of $\beta$ is presented. In our method, $\beta$ is set to 8.
Due to the vector flow modification, the separate patches are only partially aligned. Thus, the seam cutting is still required to composite these patches. Note that the local seam $s_{i}$ in the aligned patches must be connected with the initial seam $S_{0}$ end to end. Otherwise, artifacts may occur at the junction of the two seams. To prevent introducing new artifacts, we constrain the local seam to connect with the initial seam by finding a label $l^{i}$ in which the data penalty for pixel $p_{j}\in\mathbf{P}_{0}^{i}$ is
  • $D_{p_{j}}(0)=0,D_{p_{j}}(1)=\infty$, if $p_{j}\in\partial\mathbf{P}_{0}^{i}$ with $l_{p_{j}}=0$ in the initial seam $S_{0}$;
  • $D_{p_{j}}(0)=\infty,D_{p_{j}}(1)=0$, if $p_{j}\in\partial\mathbf{P}_{0}^{i}$ with $l_{p_{j}}=1$ in the initial seam $S_{0}$;
@@
where $\infty$ is set to avoid mislabeling, $\partial\mathbf{P}_{0}^{i}$ represents the border of $\mathbf{P}_{0}^{i}$. Fig. 3 shows an example including the seam merging process. We can see that the local seam is accurately merged with the initial seam.
We summarize our method in Algorithm 1.
0:  two aligned images $I_{0}$, $I_{1}$ and an initial seam $S_{0}$;
0:  final stitching seam $S_{*}$;
1:  evaluate the pixels along the seam via seam quality evaluation in Eq. (2);
2:  identify whether there are misaligned pixels along the seam;
@@
4:     extract misaligned components $\{\mathbf{C}_{i}\}_{i=1}^{n}$ in seam $S_{0}$ based on threshold $\tau$;
5:     for each $\mathbf{C}_{i}$ do
@@
7:        estimate the SIFT flow [34] between $\mathbf{P}^{i}_{0}$ and $\mathbf{P}^{i}_{1}$ and obtain the vector flow $V_{i}$;
8:        modify the vector flow via the sigmoid function $f$ in Eq. (5);
9:        align $\mathbf{P}^{i}_{0}$ and $\mathbf{P}^{i}_{1}$ via $f(V_{i})$;
10:        calculate a local stitching seam $s_{i}$ in the two aligned patches;
11:        obtain a new seam $S_{i}$ by merging $s_{i}$ into the last seam in the corresponding position.
12:     end for
13:  end if
14:  return  $S_{*}=S_{n}$
Algorithm 1 Local patch alignment module for large parallax image stitching.
IV Experimental evaluations
@@
In our experiments, the patch size for seam quality evaluation is set to $21\times 21$. Input images are aligned using REW [6], a state-of-the-art warping method. To comprehensively assess the proposed method, LPAM is applied to several advanced seam cutting methods, including basic Euclidian-based seam cutting (Baseline), perception-based seam cutting (Perception) [18], iterative seam cutting (Iterative) [19], quaternion-based seam cutting (Qua-ternion) [20], the final mosaic is generated by applying gradient domain fusion [35] to the seam cutting result. We test our method on three large parallax image datasets: 35 image pairs from Parallax [24], 24 image pairs from SEAGULL [25], and 20 image pairs from our dataset, as shown in Fig. 4. Our dataset is available in our code [^1] https://github.com/tlliao/LPAM_seam-cutting.
IV-A Quantitative evaluation

### v3

@@ Abstract
Seam cutting has shown significant effectiveness in the composition phase of image stitching, particularly for scenarios involving parallax. However, conventional implementations typically position seam-cutting as a downstream process contingent upon successful image alignment. This approach inherently assumes the existence of locally aligned regions where visually plausible seams can be established. Current alignment methods frequently fail to satisfy this prerequisite in large parallax scenarios despite considerable research efforts dedicated to improving alignment accuracy. In this paper, we propose an alignment-compensation paradigm that dissociates seam quality from initial alignment accuracy by integrating a Local Patch Alignment Module (LPAM) into the seam-cutting pipeline. Concretely, given the aligned images with an estimated initial seam, our method first identifies low-quality pixels along the seam through a seam quality assessment, then performs localized SIFT-flow alignment on the critical patches enclosing these pixels. Finally, we recomposite the aligned patches using adaptive seam-cutting and merge them into the original aligned images to generate the final mosaic. Comprehensive experiments on large parallax stitching datasets demonstrate that LPAM significantly enhances stitching quality while maintaining computational efficiency. The code is available at https://github.com/tlliao/LPAM_seam-cutting.
1 Introduction
Image stitching, a cornerstone technology in computer vision with decades of development [27], fundamentally comprises two critical phases: image alignment [4, 9, 30, 16, 11, 17, 36, 14, 18, 34, 33, 5], composition and image blending [13, 32, 12, 19, 20, 35]. Despite remarkable progress, parallax artifacts persist as the Achilles heel in consumer-level photography due to unconstrained capture conditions. Current alignment frameworks frequently fail to address this challenge adequately, making the composition stage ultimately decisive for output quality.
Contemporary solutions predominantly employ two strategies: seam-cutting [19, 12, 20, 7, 15, 35] identifying optimal transition boundaries, and seam-driven stitching [10, 31, 22] choosing an optimal seam from multiple alignment candidates. However, these post-alignment approaches inherit fundamental limitations: their effectiveness strictly depends on initial alignment accuracy. As shown in Fig. 1(b), significant parallax causes local misalignment that conventional seam-cutting methods cannot rectify, leading to visually disruptive artifacts (Fig. 1(c,d)).
@@ 1 Introduction
Figure 1: Issues of the state-of-the-art seam-cutting methods for large parallax image stitching (marked by red arrows). (b) Aligned images are obtained via method [16], where the overlapping region is averaged. The curbs are notably misaligned (Best view in color and zoom in).
@@
Our key insight is that localized post hocc alignment augmentation can overcome this bottleneck. We propose the Local Patch Alignment Module (LPAM), a lightweight add-on for existing seam-cutting pipelines. Specifically, given aligned images with an initial seam, we first identify low-quality pixels along the seam through a seam quality metric. Then, we separate their enclosing patches in the aligned images and perform localized SIFT-flow alignment for the patches. Finally, we recomposite the aligned patches using adaptive seam-cutting and merge them into the original aligned images to generate the final mosaic. We conduct extensive experiments to validate that adding LPAM into a seam-cutting can effectively and efficiently improve the stitching results, as shown in Fig. 1(e). Our core contributions are threefold:
  • A novel alignment-compensation paradigm that dissociates seam quality from initial alignment accuracy, overcoming a fundamental limitation in current methodologies
  • The first modular framework enabling seamless integration of local alignment refinement into arbitrary seam-cutting pipelines.
  • An effective implementation achieving stitching quality improvement without requiring architectural changes to host algorithms.
2 Related Work
@@
2.1 Seam-cutting methods
@@
The seam-cutting approach is a powerful and widely used composition method [13, 2, 32, 12, 19, 20, 24], which is usually expressed in terms of energy minimization and minimized via graph-cut optimization [3]. Agarwala et al. [2] applied seam-cutting to generate a “digital photomontage”. They defined the energy function as the sum of a data penalty selected by the user and an interaction penalty involving the colors, gradients, and edges of an image.
@@
Zhang et al. [32] combined the alignment error and color difference in the energy function to find a better seam that avoids splitting misaligned pixels with similar colors. Li and Liao [19] considered the non-linearity and non-uniformity of human perception into seam-cutting and defined a sigmoid-metric color difference and saliency weights in the energy function. To avoid objects from being cropped, omitted, or duplicated, Herrmann et al. [12] used object detection to match objects between images and introduced three new terms to the traditional energy function that address the above issues, respectively. Liao et al. [20] indicated that the seam with minimal cost is not necessarily optimal in terms of human perception. Then, they proposed an iterative seam estimation method where a quality evaluation algorithm for pixels along the seam is designed to guide the iteration procedure. Dai et al. [7] proposed the first end-to-end deep learning framework for image composition. They cast the seam-cutting as an image-blending problem and designed an edge-guided network to regress the blending weights and seamlessly produce the stitched image. Li and Zhou [15] utilize the quaternion representation of the color image to define the energy function and perform seam-cutting completely in the quaternion domain.
2.2 Seam-driven methods
Figure 2: An overview of the proposed local patch alignment module. Given two aligned images, we first estimate an initial seam and evaluate the pixels’ quality along the seam. Then, we adopt two stages to address the low-quality pixels: (1) Extract their enclosing patches in the aligned images for further realignment. (2) Realign each two patches using localized SIFT-flow, composite the aligned patches via a local seam-cutting.
@@ 2.2 Seam-driven methods
The seam-driven strategy was first proposed by Gao et al. [10]. They indicated that the perceptually best result is not necessarily from the best global alignment. Thus, they generated finite alignment candidates from multiple RANSAC [8] procedures and applied seam-cutting to estimate multiple seams from these candidates. Then, a seam quality metric is defined to evaluate these seams, and the seam with minimal cost is chosen to produce a final mosaic. Zhang and Liu [31] improved the screening process of alignment candidates by combining homography with content-preserving warps. They used the seam cost of the energy function to evaluate the seam quality. Lin et al. [22] generated the alignment candidates via a superpixel-based feature grouping and a structure-preserving warp. The warp is gradually improved via adaptive feature weighting. They also defined a seam quality metric based on the zero-mean normalized cross-correlation (ZNCC) score. Zhang et al. [35] proposed an iterative seam-based stitching method that leverages dense flow estimation to directly estimate the seam in the input images and use it to solve a spatial smooth warp.
@@
Both seam-cutting and seam-driven methods focus on defining various energy functions to find an optimal seam from the settled pre-aligned images. For images with large parallax, these methods may easily fail if the alignment methods fail to align the images locally. Our method, on the other hand, improves performance by realigning the locally misaligned patches where the estimated seam passes.
3 Proposed Method
@@
As depicted in Fig. 2, our method consists of two main stages, seam quality evaluation and patch extraction, succeeded by patch alignment and seam-merging. Subsequently, we first provide a preliminary of the basic seam-cutting algorithm and then present a detailed description of the two stages.
3.1 Preliminary
@@
Given aligned target image $I_{0}$ and reference image $I_{1}$, for the overlapping region $\mathbf{O}$, seam-cutting method estimates an initial seam $S_{0}$ by finding a labeling $l$ that minimizes the energy function
@@
3.2 Seam quality evaluation and patch extraction
Since the seam-cutting methods mainly fail due to the structure misalignment, we use the structural similarity (SSIM) index [28] to evaluate each pixel $p_{i}$ on the seam $S_{0}$. Concretely, we extract two local patches $P_{0}$ and $P_{1}$ centered at $p_{i}$, and define the evaluation error for $p_{i}$ as
@@ 3.2 Seam quality evaluation and patch extraction
Figure 3: Illustration for patch alignment and seam-merging process. Artifacts are denoted by red arrows. From left to right in (a,b,c,d): vector flow, local stitching seam, and local composited result. For clarity, only part of the vector flow is presented. From (a) to (d): results of original $V^{i}$, modified $f(V^{i})$ with $\beta=1,4,8$ in Eq. (5), respectively. The green rectangle in the middle of each subfigure indicates the position of the aligned patch and the merging position of the local seam. The right image in each subfigure shows the composited patch result via the seam-cutting in the middle (Best view in color and zoom in).
@@
We set $k$ to 1.5 in our method. The threshold $\tau$ divides the seam into multiple components. Notably, the threshold might not be sufficiently precise to identify all misaligned pixels. However, we allow for some false positives because the subsequent alignment process does not yield inferior results. For each misaligned component $\mathbf{C}_{i}$, we separate its enclosing rectangular patches $\mathbf{P}^{i}_{0}$ and $\mathbf{P}^{i}_{1}$ from the corresponding aligned images (see Fig. 2).
3.3 Patch alignment and seam-merging
3.3.1 Patch alignment
The rectangular patches manifest structural misalignment stemming from the feature-based alignment algorithm, primarily due to insufficient feature correspondences in texture-deficient regions. Thus, we utilize SIFT flow [23] to construct dense correspondences among the pixels within the patches. Theoretically, for each pixel $p_{j}$ in $\mathbf{P}_{0}^{i}$, there exists a shift vector $V^{i}(p_{j})$ such that
@@ 3.3.1 Patch alignment
Then, the two patches can be aligned via the vector flow $V^{i}$. However, if the boundary pixels of the patch $\mathbf{P}_{0}^{i}$ have large shift vectors, merging the aligned patch into the initial aligned result may introduce more artifacts around the patch, as shown in Fig. 3. To address this, we smoothly modify the vector flow and realign the two patches as
$\tilde{\mathbf{P}}_{0}^{i}(p_{j})=\mathbf{P}_{0}^{i}(p_{j}+f(p_{j})\cdot V^{i}(p_{j})),\tilde{\mathbf{P}}_{1}^{i}(p_{j})=\mathbf{P}_{1}^{i}(p_{j})$ (4)
@@
For horizontal patch alignment, $t$ represents the $x$-coordinate of the pixel $p_{j}$ within patch $\mathbf{P}_{0}^{i}$, normalized to $[0,1]$. $\beta$ controls the rate of change of the function $f$. The sigmoid function ensures that the warped patch $\tilde{\mathbf{P}}_{0}^{i}$ is smoothly consistent with $I_{0}$ when $t=0$ and with $I_{1}$ when $t=1$. The warped patches $\tilde{\mathbf{P}}_{0}^{i}$ and $\tilde{\mathbf{P}}_{1}^{i}$ are then placed back to the original aligned images $I_{0}$ and $I_{1}$. Fig. 3 shows a comparison example of the modified vector flow under different settings of $\beta$. In our method, $\beta$ is set to 8.
3.3.2 Seam-merging
Due to the vector flow modification, the rectangular patches are only partially aligned. Thus, seam-cutting is still required to composite these patches. Note that the local seam $s_{i}$ in the aligned patch must be connected with the initial seam $S_{0}$ end to end. Otherwise, artifacts may occur at the junction of the two seams. To prevent introducing new artifacts, we constrain the local seam to connect with the initial seam by finding a label $l^{i}$ in which the data penalty for pixel $p_{j}\in\tilde{\mathbf{P}}_{0}^{i}$ is
  • $D_{p_{j}}(0)=0,D_{p_{j}}(1)=\infty$, if $p_{j}\in\partial\tilde{\mathbf{P}}_{0}^{i}$ with $l_{p_{j}}=0$ in the initial seam $S_{0}$;
  • $D_{p_{j}}(0)=\infty,D_{p_{j}}(1)=0$, if $p_{j}\in\partial\tilde{\mathbf{P}}_{0}^{i}$ with $l_{p_{j}}=1$ in the initial seam $S_{0}$;
@@ 3.3.2 Seam-merging
where $\infty$ is set to avoid mislabeling, $\partial\tilde{\mathbf{P}}_{0}^{i}$ represents the border of the warped patch $\tilde{\mathbf{P}}_{0}^{i}$. Fig. 3 shows an example including the seam-merging process. We can see that the local seam in the aligned patch is accurately merged with the initial seam. After realigning all the rectangular patches, we can generate the final mosaic by compositing the locally realigned images $\tilde{I}_{0}$ and $\tilde{I}_{1}$ with the new stitching seam. We summarize our method in Algorithm 1.
Algorithm 1 Local patch alignment module for large parallax image stitching
0:  aligned images $I^{0}_{0}$, $I^{0}_{1}$ and an initial seam $S_{0}$
0:  final seam $S_{*}$ and realigned images $I^{*}_{0}$, $I^{*}_{1}$
1:  evaluate $S_{0}$ via Eq. (2);
2:  identify whether there are misaligned pixels along $S_{0}$;
@@
4:     extract misaligned components $\{\mathbf{C}_{i}\}_{i=1}^{n}$ in $S_{0}$;
5:     for $i=1$ todo
@@
7:        obtain $V_{i}$ via Eq. (3);
8:        modify the vector flow via $f$ in Eq. (5);
9:        obtain $\tilde{\mathbf{P}}_{0}^{i}$ and $\tilde{\mathbf{P}}_{1}^{i}$ via Eq. (4);
10:        calculate a local stitching seam $s_{i}$ in $\tilde{\mathbf{P}}_{0}^{i}$ and $\tilde{\mathbf{P}}_{1}^{i}$;
11:        $S_{i}:=S_{i-1}\leftarrow s_{i}$;
12:        $I^{i}_{0}:=I^{i-1}_{0}\leftarrow\tilde{\mathbf{P}}_{0}^{i}$;
13:        $I^{i}_{1}:=I^{i-1}_{1}\leftarrow\tilde{\mathbf{P}}_{1}^{i}$.
14:     end for
15:  end if
16:  return  $S_{*}=S_{n},I^{*}_{0}=I^{n}_{0},I^{*}_{1}=I^{n}_{1}$
@@
4 Experiments
Figure 5: Visual seam quality comparisons w/o and w/ LPAM. Top: initial seams w/o LPAM. Bottom: final seams w/ LPAM.
In the experiments, the patch size for seam quality evaluation is set to $21\times 21$. Input images are aligned using REW [16], a state-of-the-art warping method. To comprehensively assess the proposed method, LPAM is applied to several advanced seam-cutting methods, including basic Euclidian-based seam-cutting (Baseline), perception-based seam-cutting (Perception) [19], iterative seam-cutting (Iterative) [20], and quaternion-based seam-cutting (Quaternion) [15], the final mosaic is generated by simply compositing the aligned images according to the stitching seam. We test our method on three large parallax image stitching datasets: 35 image pairs from Parallax [31], 24 image pairs from SEAGULL [22], and 20 image pairs from our dataset, as shown in Fig. 4.
Table 1: Quantitative evaluation of applying LPAM to different seam-cutting methods under the same alignment results via REW [16].
Dataset Parallax [31] SEAGULL [22] Our dataset
+LPAM 0.078 31.85 0.782 0.100 0.096 28.84 0.697 0.137 0.075 28.76 0.778 0.104
+LPAM 0.074 31.23 0.795 0.084 0.103 25.37 0.682 0.127 0.072 28.82 0.801 0.084
+LPAM 0.076 30.87 0.800 0.087 0.103 25.37 0.682 0.127 0.072 28.91 0.786 0.101
+LPAM 0.080 29.41 0.783 0.085 0.105 26.09 0.677 0.136 0.067 29.25 0.812 0.078
Table 2: Quantitative evaluation of applying LPAM to basic Euclidian-based seam-cutting (Baseline) under different alignment results.
Dataset Parallax [31] SEAGULL [22] Our dataset
Homo 0.103 28.17 0.692 0.157 0.123 25.79 0.622 0.192 0.104 25.71 0.706 0.134
+LPAM 0.092 28.78 0.725 0.133 0.107 28.50 0.673 0.165 0.094 26.22 0.742 0.114
ANAP [21] 0.091 27.99 0.718 0.150 0.105 26.07 0.669 0.183 0.087 27.11 0.704 0.137
+LPAM 0.081 28.64 0.766 0.123 0.093 26.88 0.726 0.146 0.080 27.73 0.749 0.112
GSP [6] 0.078 30.14 0.762 0.105 0.102 26.34 0.670 0.161 0.081 28.36 0.746 0.117
+LPAM 0.072 30.65 0.798 0.088 0.092 29.53 0.719 0.133 0.077 28.74 0.768 0.108
REW [16] 0.088 31.01 0.750 0.119 0.106 26.20 0.638 0.163 0.079 28.50 0.756 0.114
+LPAM 0.078 31.85 0.782 0.100 0.096 28.84 0.697 0.137 0.075 28.76 0.778 0.104
UDIS++ [25] 0.120 27.05 0.643 0.189 0.120 25.58 0.589 0.212 0.099 26.34 0.658 0.164
+LPAM 0.100 28.33 0.705 0.149 0.109 26.32 0.651 0.174 0.088 27.04 0.715 0.130
4.1 Quantitative evaluation